# Large Language Models in Argument Mining: A Survey

### v5

@@ 5.3 Evolution of Evaluation Metrics
Early corpora with nominal “good / bad” labels reported Cohen’s (Hsu and Field, 2003) or Fleiss’ $\kappa$ (Fleiss et al., 1981). As datasets grew and adopted ordinal or continuous scales, Krippendorff’s $\alpha$ (Krippendorff, 2011) became preferred because it handles mixed measurement levels (Stab and Gurevych, 2017). Large crowdsourced resources such as IBM Rank-30k now accompany raw labels with MACE-style worker weighting; after filtering, $\alpha$ values can reach 0.93, approaching expert reliability (Gretz et al., 2020). A recent meta-survey corroborates these gains across 32 datasets (Ivanova et al., 2024).
@@
When AQA was framed as binary classification, macro-averaged Precision, Recall, and F1 sufficed (Wachsmuth et al., 2017a). As labels became graded, regression metrics—mean absolute error and root MSE—and rank correlations (Pearson (Benesty et al., 2009), Spearman $\rho$ (Hauke and Kossowski, 2011)) took centre stage. Persing & Ng first cast persuasiveness as a six-point regression task (Persing and Ng, 2020), and Lauscher et al. popularised Spearman $\rho$ on GAQCorpus (Lauscher et al., 2020). Multi-task heads that jointly predict stance and several quality facets are usually evaluated with a macro-averaged MAE or F1 over dimensions (Favreau et al., 2022).
@@
[…]
@@
Recent studies measure AQA inside larger pipelines. The COVERAGE metric scores how thoroughly a summary reflects argumentative key points without gold labels, serving as an unsupervised proxy for sufficiency (Khosravani et al., 2024). In the AMELIA benchmark for Italian law, system impact is quantified by the rise in judge agreement when automatically ranked arguments are added to briefs (Grundler et al., 2024). Soft-label evaluation—comparing predicted distributions to annotator distributions—operationalises Plank’s call to treat disagreement as signal (Plank, 2022).
@@ 5.4 Applications and Future Works
Search engines that index argumentative content (e.g. args.me) already re-rank results by predicted cogency, sufficiency or style (Wachsmuth et al., 2017b). IBM Debater uses pairwise convincingness scores to surface the strongest counter-points in debate preparation (Gretz et al., 2020). Empirical studies show that integrating AQA scores into BM25 or dense-retrieval pipelines raises NDCG@555 without additional user feedback (Gienapp et al., 2020).
@@
Automated writing-evaluation tools such as ArgRewrite provide sentence-level suggestions that target logical sufficiency and rhetorical clarity; classroom deployments report statistically significant gains in argumentative essay scores (Favreau et al., 2022). In science-education contexts, mapping Walton’s Critical Questions to teacher feedback helps novices revise reasoning chains more effectively (Winkle et al., 2020).
@@
The legal benchmark AMELIA attaches sufficiency and statute-relevance scores to Italian court opinions, enabling intelligent brief-generation and judge-support systems (Grundler et al., 2024). In corporate governance, argument-mapping dashboards highlight reasoning gaps in environmental, social and governance (ESG) reports, assisting compliance officers (Palmieri, 2024).
@@ 6 Modelling Paradigms in the LLM Era
Large Language Models (LLMs) have transformed argument mining from a pipeline of narrow, task-specific classifiers into a set of interaction paradigms in which one or more foundation models collaborate with retrieval modules, symbolic structures or lightweight adapters. Rather than ask whether an architecture is “better” than another, current research asks how LLMs are enlisted as zero-shot predictors, few-shot imitators, knowledge retrievers, judges, data generators, or all of the above. This section synthesises the dominant paradigms, emphasising the design rationales, empirical pay-offs, and unresolved pitfalls.
@@ 6.1 Prompt-based Argument Mining
Early demonstrations showed that even naïve declarative prompts (“Is the following sentence a claim?”) enable GPT-3 to match feature-rich baselines on component segmentation and stance detection without gradient updates (Chen et al., 2024). Subsequent work moved beyond template engineering to embed task structure directly inside the prompt. Altemeyer et al. (2025) inserted chain-of-thought (CoT) exemplars that model the logical leap between premise and conclusion; the reasoning trace, rather than the final label, accounted for most of the performance gain on sufficiency and convincingness. Prompting can also express task hierarchy: the PITA framework conditions relation prediction on previously elicited component spans within the same conversation context and reports significant improvements on Reddit and legal briefs (Sun et al., 2024b). Role prompts further reduce topical and demographic bias: framing the model as a Debate Judge or Devil’s Advocate attenuates polarity skew in political fact-checking (Dayanik and Padó, 2020).
Yet prompt brittleness persists. Small lexical tweaks can flip stance labels, and CoT traces lengthen latency in interactive writing tools. Robust prompt ensembles and self-consistency voting help but introduce computational overhead; empirical studies suggest latency grows almost linearly with ensemble size (Zhang et al., 2025). Designing prompts that capture argumentation theory while remaining compact therefore remains an open problem.
6.2 Parameter-efficient Instruction and Multi-task Tuning
Zero-shot prompting rarely suffices for domain-sensitive genres such as biomedical abstracts or statutory law. Instruction tuning—fine-tuning on a mixture of task-specific exemplars framed as natural-language instructions—bridges much of this gap. A single T5-base model (Chung et al., 2024) jointly instruct-tuned on claim segmentation, relation prediction and quality scoring yields consistent gains across all three tasks on IAM and ARIES, while adapters keep GPU memory within budget (Cheng et al., 2022; Gemechu et al., 2024). In the legal domain, ArgLegalSumm couples instruction tuning with an argument salience re-ranker, improving summary coherence in Italian case law (Elaraby et al., 2023). Cross-domain continual learning is more fragile: when a biomedical corpus is appended late in the schedule, earlier gains on social-media stance detection erode, echoing catastrophic forgetting observed in multi-task translation (Pang et al., 2024).
Recent work therefore explores parameter-efficient updates: LoRA (Hu et al., 2022), prefix-tuning and bit-fit adapters enable domain adaptation with two orders of magnitude fewer trainable parameters while preserving general AM capability (Li et al., 2024b). Still unsettled is how to balance tasks of differing label density: curriculum strategies that start with coarse component labels before introducing noisy quality scores appear promising, but systematic ablation is scarce.
@@ 6.3 Retrieval-Augmented Generation and Reasoning
LLMs remain constrained by context windows and training cut-off dates. Retrieval-augmented generation (RAG) mitigates both limits by injecting external evidence into the prompt. In debate summarisation, RAG systems retrieve topically relevant counter-arguments and background facts, reducing hallucination and increasing factual density in abstractive summaries (Syed et al., 2023). The OpenDebateEvidence benchmark pairs fifteen million statements with evidence links, enabling end-to-end retrieval-generation pipelines that outperform text-only baselines by a meaningful margin (Roush et al., 2024). Retrieval can also be argument-theoretic: semantic-graph matching over Kialo debate trees yields higher recall of undercutters and mitigations than purely lexical search (Ruiz-Dolz et al., 2023).
Several challenges loom. Evidence ranking often inherits popularity biases, favouring mainstream sources over minority viewpoints. Long evidence snippets may eclipse the claim inside the prompt, forcing heuristic truncation. Finally, retrieval latency threatens interactive applications; hybrid dense-sparse indexing and passage caching alleviate but do not solve the issue.
@@ 6.4 Synthetic Supervision and Data Augmentation
LLMs themselves can generate training data, especially where expert annotation is scarce. Wagner et al. (2025) show that LLM-synthesised stance corpora, when filtered for diversity, raise low-resource performance more than in-context few-shot learning. Similarly, LLM-generated key-point coverage scores serve as weak labels for automatic evaluation metrics, correlating well with human judgements (Khosravani et al., 2024). Synthetic supervision also facilitates controlled experiments: topic-balanced sampling can equalise stance priors across political entities, supporting fairness audits (Mehrabi et al., 2022).
The downside is distributional drift. Synthetic arguments tend to be lexically fluent, hedge-free and uniformly polite, properties that can over-regularise fine-tuned models. Provenance cards documenting prompt templates, sampling temperatures, rejection rates and post-editing guidelines are therefore being proposed as a best practice . Legal uncertainty over copyright of machine-generated text remains unresolved.
@@ 7 Future Directions
As argument mining transitions into the era of large language models (LLMs), a wide range of opportunities emerge alongside challenges that must be systematically addressed. This chapter synthesizes key limitations identified across prior sections and outlines promising future research directions spanning technical, theoretical, and practical dimensions.
@@ 7.1 Interpretability and Controllability
LLM outputs can be impressively fluent yet logically opaque. Black-box convincingness scores or abstractive summaries are difficult to trust in law, medicine or policy making, where explanations must be auditable and contestable. Promising directions include symbolic scaffolding, in which an LLM populates a Toulmin-style or Walton-style template before emitting a verdict (Lee et al., 2024; Borg and Bex, 2024); rationale-aligned distillation, where model predictions are accepted only if an intermediate reasoning trace satisfies argument-quality criteria (Zhao et al., 2024); and rule-augmented decoding, which filters candidate continuations against critical-question constraints during generation (Palmieri, 2024). Empirical studies show that such structured intermediates improve user trust more than generic saliency maps, yet rigorous human-in-the-loop evaluations are still rare.
7.2 Data Quality, Bias and Hallucination
LLM training corpora contain ideological skew, demographic imbalance and source-credibility noise. These artefacts surface in persuasive writing: emotionally charged but weakly supported arguments are often overrated by LLM judges (Chen and Eger, 2025). Counterfactual calibration and actor masking reduce partisan bias in stance detection (Li et al., 2025a; Dayanik and Padó, 2020), but systematic audits across domains remain scarce. Hallucination is the sister problem: retrieval-augmented generation mitigates unsupported claims yet inherits popularity and language biases from search engines (Roush et al., 2024). Future benchmarks should combine bias slices e.g. position papers from under-represented groups— with factual stress tests to quantify both issues jointly.
@@ 7.3 Generalisation and Robustness
Most LLM studies report single-domain results; cross-genre transfers often collapse. Debate summarisers tuned on English legal cases lose more than a third of their content fidelity when applied to multi-party chat (Zhao et al., 2023). Adversarial paraphrase attacks flip stance predictions with minimal lexical edits (Gienapp et al., 2020). Potential remedies include curriculum-style multi-task fine-tuning, meta-learning over domain clusters, and evaluation suites that mix genres, languages and modalities within the same test fold.
@@ 7.4 Synthetic Supervision and Benchmark Integrity
LLM-generated corpora such as ZeroStance (Zhao et al., 2024) and OpenDebateEvidence (Roush et al., 2024) accelerate data growth but raise three red flags: (i) contamination models may re-encounter their own pre-training shards; (ii) superficial reasoning synthetic counter-claims are stylistically rich yet logically shallow; (iii) blind spot edge cases rarely appear in temperature-controlled sampling. Hybrid pipelines that blend human editing, diversity-aware sampling and automatic quality-metrics (e.g. coverage and exhaustiveness scores from Khosravani et al. (2024)) are a promising compromise. Benchmark leaders should publish provenance cards documenting prompt templates, sampling temperatures and rejection rates.

### v6

@@ 5.3 Evolution of Evaluation Metrics
Early corpora with nominal “good / bad” labels reported Cohen’s (hsu2003interrater) or Fleiss’ $\kappa$ (fleiss1981measurement). As datasets grew and adopted ordinal or continuous scales, Krippendorff’s $\alpha$ (krippendorff2011computing) became preferred because it handles mixed measurement levels (DBLP:conf/eacl/GurevychS17). Large crowdsourced resources such as IBM Rank-30k now accompany raw labels with MACE-style worker weighting; after filtering, $\alpha$ values can reach 0.93, approaching expert reliability (DBLP:conf/aaai/GretzFCTLAS20). A recent meta-survey corroborates these gains across 32 datasets (DBLP:conf/emnlp/IvanovaHN24).
@@
When AQA was framed as binary classification, macro-averaged Precision, Recall, and F1 sufficed (DBLP:conf/eacl/WachsmuthSHPBHN17). As labels became graded, regression metrics—mean absolute error and root MSE—and rank correlations (Pearson (benesty2009pearson), Spearman $\rho$ (hauke2011comparison)) took centre stage. Persing & Ng first cast persuasiveness as a six-point regression task (DBLP:conf/lrec/PersingN20), and Lauscher et al. popularised Spearman $\rho$ on GAQCorpus (DBLP:conf/coling/LauscherNNT20). Multi-task heads that jointly predict stance and several quality facets are usually evaluated with a macro-averaged MAE or F1 over dimensions (DBLP:conf/flairs/FavreauZB22).
@@
For systems that must order arguments, pairwise and listwise metrics dominate. Pairwise Accuracy (PAcc) on ConvArg and IBM-Pairs remains the primary measure of relative convincingness (DBLP:conf/acl/HabernalG16). Retrieval-oriented work adds MRR and NDCG@$k$; PageRank-style relevance improved NDCG by 4–5 points and is evaluated with Spearman $\rho$ against human rankings (DBLP:conf/eacl/WachsmuthSA17). Efficient pairwise annotation plus ranking loss further increases Kendall $\tau$ while halving labelling cost (DBLP:conf/acl/GienappSHP20).
@@
Recent studies measure AQA inside larger pipelines. The COVERAGE metric scores how thoroughly a summary reflects argumentative key points without gold labels, serving as an unsupervised proxy for sufficiency (DBLP:conf/naacl/Khosravani0T24). In the AMELIA benchmark for Italian law, system impact is quantified by the rise in judge agreement when automatically ranked arguments are added to briefs (DBLP:conf/clic-it/GrundlerGSFGPLS24). Soft-label evaluation—comparing predicted distributions to annotator distributions—operationalises Plank’s call to treat disagreement as signal (DBLP:conf/emnlp/Plank22).
@@ 5.4 Applications and Future Works
Search engines that index argumentative content (e.g. args.me) already re-rank results by predicted cogency, sufficiency or style (DBLP:conf/eacl/WachsmuthSA17). IBM Debater uses pairwise convincingness scores to surface the strongest counter-points in debate preparation (DBLP:conf/aaai/GretzFCTLAS20). Empirical studies show that integrating AQA scores into BM25 or dense-retrieval pipelines raises NDCG@$5$ without additional user feedback (DBLP:conf/acl/GienappSHP20).
@@
Automated writing-evaluation tools such as ArgRewrite provide sentence-level suggestions that target logical sufficiency and rhetorical clarity; classroom deployments report statistically significant gains in argumentative essay scores (DBLP:conf/flairs/FavreauZB22). In science-education contexts, mapping Walton’s Critical Questions to teacher feedback helps novices revise reasoning chains more effectively (DBLP:conf/icls/WinklePN20).
@@
The legal benchmark AMELIA attaches sufficiency and statute-relevance scores to Italian court opinions, enabling intelligent brief-generation and judge-support systems (DBLP:conf/clic-it/GrundlerGSFGPLS24). In corporate governance, argument-mapping dashboards highlight reasoning gaps in environmental, social and governance (ESG) reports, assisting compliance officers (DBLP:conf/cmna/Palmieri24).
@@ 6 Modelling Paradigms in the LLM Era
Large Language Models (LLMs) have reshaped argument mining (AM) from a pipeline of specialised classifiers into a set of model–interaction paradigms. Rather than asking which architecture performs best, current research examines how foundation models are enlisted—as zero-shot predictors, few-shot mimics, retrieval-aware reasoners, synthetic data generators, or automated judges—and how these roles alter the design of AM tasks and evaluation. This section consolidates four dominant paradigms: prompting, instruction-based multi-task tuning, retrieval-augmented reasoning, and synthetic supervision. Across these paradigms, we highlight their methodological motivations, empirical benefits, and the pitfalls that motivate redesigned evaluation practices.
@@ 6.1 Prompt-based Argument Mining
Prompting constitutes the most lightweight and arguably the most transformative entry point for LLM use in AM. Even simple declarative prompts (e.g., “Is the following sentence a claim?”) enable GPT-style models to match classical feature-rich baselines on claim segmentation and stance detection without gradient updates (DBLP:conf/acl/ChenCLB24). Subsequent work moves beyond surface templates to leverage explicit argumentative structure inside the prompt. DBLP:journals/corr/abs-2503-00847 show that inserting chain-of-thought exemplars produces more faithful reasoning traces for sufficiency and convincingness assessment; the intermediate rationale, rather than the label, accounts for most of the gain. Hierarchical prompting strategies such as PITA condition relation prediction on prior component segmentation within the same conversation context (DBLP:conf/acl/SunWBL0Y0X24). Role prompts (“Debate Judge”, “Devil’s Advocate”) can attenuate partisan or demographic biases in political domains (DBLP:conf/acl/DayanikP20).
Despite its flexibility, prompting remains brittle. Small lexical perturbations can flip stance labels, and longer chain-of-thought traces incur noticeable latency in interactive systems. Self-consistency ensembles and voting can stabilise predictions but scale inference cost nearly linearly in ensemble size (DBLP:journals/corr/abs-2503-10881). Designing compact prompts that faithfully encode argumentative theory remains an open challenge.
6.2 Instruction and Multi-task Tuning
Prompting alone is insufficient for domain-sensitive genres such as biomedical abstracts or statutory law. Instruction tuning—fine-tuning on curated exemplars framed as natural-language instructions—enables a single model to accommodate multiple AM sub-tasks. For example, a T5-base model jointly tuned on claim segmentation, relation prediction, and quality scoring achieves consistent improvements across IAM and ARIES (DBLP:conf/acl/ChengBHYZS22; gemechu2024aries), while keeping memory demands modest through adapter-based updates. In legal summarisation, ArgLegalSumm couples instruction tuning with salience-aware reranking to improve coherence for Italian case law (DBLP:conf/acl/ElarabyZL23).
However, multi-task instruction tuning introduces new tensions. Adding late-stage biomedical data often erodes earlier gains on social-media stance detection, echoing catastrophic forgetting in multi-domain translation (pang2024rethinking). Parameter-efficient tuning—LoRA (DBLP:conf/iclr/HuSWALWWC22), prefix-tuning, and bit-fit adapters—offers a partial remedy, enabling domain adaptation with orders-of-magnitude fewer trainable parameters while preserving general AM capabilities (DBLP:conf/acl/0074WSBMZZWHLN24). Yet, how to balance tasks with uneven label density (e.g., frequent component spans vs. sparse quality scores) remains under-explored; preliminary curriculum approaches yield promise but lack systematic ablation.
@@ 6.3 Retrieval-Augmented Generation and Reasoning
LLMs remain constrained by fixed context windows and outdated pretraining corpora. Retrieval-Augmented Generation (RAG) mitigates both issues by injecting external evidence during inference. In debate summarisation, retrieving counter-arguments or factual background reduces hallucination and increases factual density (DBLP:conf/sigdial/SyedZHWP23). The OpenDebateEvidence benchmark (DBLP:conf/nips/RoushSBZMZBVS24) pairs 15 million argumentative statements with verifiable evidence links, enabling end-to-end retrieval–generation pipelines that outperform text-only baselines. Argument-theoretic retrieval using semantic-graph matching over debate trees improves recall for undercutters and mitigations beyond lexical search (DBLP:conf/emnlp/Ruiz-DolzHG23).
Nevertheless, RAG introduces its own difficulties. Evidence ranking inherits popularity biases, often marginalising minority viewpoints. Long passages can saturate prompts, forcing heuristic truncation that discards crucial argumentative steps. Moreover, high-latency retrieval complicates deployment in writing assistants, where hybrid dense–sparse indexing and caching help but cannot fully resolve delays.
@@ 6.4 Synthetic Supervision and Data Augmentation
LLMs now serve as data generators and weak supervisors. DBLP:conf/iclr/WagnerBZH25 show that LLM-synthesised stance datasets, when filtered for topical diversity, outperform few-shot prompting for low-resource languages. LLM-generated key-point coverage scores correlate with human judgments and can supervise automatic evaluation metrics (DBLP:conf/naacl/Khosravani0T24). Synthetic data also facilitates fairness investigations by equalising stance priors across sensitive groups (DBLP:journals/csur/MehrabiMSLG21).
Yet synthetic supervision risks distributional drift. Machine-generated arguments tend to be stylistically homogeneous—overly polite, hedge-free, lexically fluent—and may over-regularise fine-tuned models. Researchers increasingly call for dataset “provenance cards” documenting prompt templates, sampling temperatures, rejection statistics, and human post-editing procedures, especially as copyright and licensing questions remain unsettled.
6.5 Discussion: Capabilities, Pitfalls, and Open Problems
Across the four paradigms, a common pattern emerges: LLMs increase coverage and flexibility but simultaneously introduce new evaluation challenges. Prompting enables broad generalisation but is brittle; instruction tuning captures domain regularities but risks forgetting; RAG reduces hallucination but amplifies evidence-selection biases; synthetic supervision expands datasets but propagates stylistic artefacts from the generator.
These shifts complicate the task of evaluating AM systems. Classical metrics—F1, accuracy, pairwise convincingness—assume a stable ontology of components, relations, and evidence. In contrast, LLM-era systems generate rationales, retrieve external evidence, or produce synthetic training samples, blurring the boundary between “task performance” and “data production”. Evaluation therefore becomes not only a matter of correctness but also of robustness, provenance, bias auditing, and consistency across modalities and domains. As discussed in Chapter 5, modern evaluation must treat AM systems as multi-component agents rather than isolated classifiers.
Open challenges remain substantial: (1) designing prompt-robust, reproducible evaluation protocols; (2) developing multi-task curricula that preserve earlier capabilities while adding new ones; (3) creating retrieval benchmarks that reflect viewpoint diversity rather than popularity; (4) establishing principled standards for synthetic data documentation and bias auditing. Addressing these problems will determine whether LLM-enhanced AM evolves into a reliable scientific discipline or remains a collection of loosely connected heuristics.
@@ 7 Future Directions
As argument mining enters the era of large language models (LLMs), the field faces an inflection point. The preceding chapters have shown that advances in modelling, prompting, dataset construction, and evaluation co-evolve in ways that amplify both opportunities and risks. This section synthesises the cross-cutting limitations identified throughout the survey and outlines several research trajectories that may shape the next decade of argument–mining research.
@@ 7.1 Interpretability and Controllability
LLM-based argument mining systems produce outputs that are often fluent, stylistically sophisticated, and broadly aligned with human reasoning—yet the underlying decision mechanisms remain opaque. In high-stakes settings such as law, medicine, or policy analysis, opaque convincingness scores or abstractive summaries are difficult to verify or contest. Recent work proposes structured intermediates as a bridge between symbolic argumentation theory and neural generation. Examples include filling Toulmin- or Walton-style templates prior to issuing a verdict (DBLP:journals/peerj-cs/LeeLK24; DBLP:journals/ijar/BorgB24), rationale-constrained distillation of LLM judgments (DBLP:conf/acl/ZhaoLCZ24), and rule-augmented decoding that checks candidate continuations against critical questions during generation (DBLP:conf/cmna/Palmieri24). Early empirical evidence suggests such scaffolding improves user trust more than generic saliency maps, yet rigorous human-in-the-loop evaluations remain rare. Developing frameworks that support contestability, error diagnosis, and user-guided correction represents a central open challenge.
7.2 Data Quality, Bias, and Hallucination
The quality of LLM-driven argument mining is inevitably bounded by the quality and diversity of its training data. Pre-training corpora encode ideological skew, demographic imbalance, and heterogeneous source credibility, all of which can surface in persuasive writing tasks: emotionally charged yet weakly supported arguments tend to be overrated by LLM judges (DBLP:journals/corr/abs-2503-00024). Existing mitigation strategies—such as actor masking (DBLP:conf/acl/DayanikP20) and counterfactual calibration for stance detection (DBLP:conf/naacl/LiZLGWZLWX25)—address slices of the problem but fall short of comprehensive bias auditing across domains and languages. Hallucination introduces a complementary failure mode. Although retrieval-augmented pipelines reduce unsupported claims, they inherit biases from search engines and ranking algorithms (DBLP:conf/nips/RoushSBZMZBVS24). Future benchmarks should therefore combine bias slices—e.g., arguments produced by under-represented communities—with factual stress tests that expose models to ambiguous, adversarial, or conflicting evidence.
@@ 7.3 Generalisation and Robustness
Despite encouraging single-domain results, robust generalisation remains elusive. Systems tuned on legal cases often experience sharp declines when applied to multi-party chat debates or open-domain social media (DBLP:conf/emnlp/ZhaoWP23). Stance detectors remain vulnerable to adversarial paraphrases that alter only surface form (DBLP:conf/acl/GienappSHP20). These issues highlight the fragility of learned argumentative representations and the limitations of narrow training distributions. Promising paths forward include curriculum-style multi-task tuning, meta-learning over domain clusters, and the development of evaluation suites that interleave genres, modalities, and languages within the same test fold. Robustness must also be examined beyond accuracy: future systems should report calibration under uncertainty, degradation curves under distribution shift, and performance on deliberately challenging cases.
@@ 7.4 Synthetic Supervision and Benchmark Integrity
LLM-generated datasets—such as ZeroStance (DBLP:conf/acl/ZhaoLCZ24) and OpenDebateEvidence (DBLP:conf/nips/RoushSBZMZBVS24)—accelerate dataset growth but introduce structural risks. Three concerns are particularly acute. First, contamination: synthetic corpora may inadvertently reproduce fragments of the model’s own pre-training data. Second, superficial reasoning: many synthetic counter-claims are stylistically well-formed yet logically shallow. Third, limited coverage: temperature-controlled sampling rarely captures rare argumentative moves or culturally specific reasoning styles. Hybrid pipelines that blend human editing with diversity-aware sampling and automatic quality metrics—such as coverage and exhaustiveness scores (DBLP:conf/naacl/Khosravani0T24)—offer a promising compromise. To preserve benchmark integrity, dataset creators should release transparent provenance cards detailing prompt templates, sampling parameters, rejection criteria, and post-editing guidelines.